OpenStudy (texmechanica):

What is the x-coordinate of the solution of the system? {5x-y=-21} {x+y=-3}

OpenStudy (ryanchick):

20R1

OpenStudy (ryanchick):

and second one -3

OpenStudy (mathmale):

@ryanchick: "helping" others does not include just handing out answers. Please show your stuff by guiding TexMechanica towards finding his/her own correct answer.

OpenStudy (mathmale):

@TexMechanica: Here you are given a system of linear equations and are asked to find the solution, or at least part of it. If you can eliminate y from both equations, you'll have one equation in x alone, that can be solved for x. Methods to consider using: substitution, addition/subtraction, graphing.

OpenStudy (ryanchick):

okay sorry im new

OpenStudy (texmechanica):

I don't even know where to begin.

OpenStudy (ryanchick):

dont worry

OpenStudy (ryanchick):

ill try and help

OpenStudy (mathmale):

{5x-y=-21} {x+y=-3} might be a little easier to visualize if you remove the brackets: 5x-y=-21 x +y=-3 Tex: simply add these two equations together, column by column. Write your result below. This is the "addition/subtraction" method of elimination.

OpenStudy (ryanchick):

@TexMechanica

OpenStudy (texmechanica):

I don't know.

OpenStudy (texmechanica):

Thanks for the help guys but i gotta go anyway.